OpenStudy (anonymous):

Please help Q2: equation of straight lines through a fixed point (h,k)..

OpenStudy (anonymous):

You need at least two points to define a line. Do you mean passing through (h,k) and the origin?

OpenStudy (anonymous):

um, I don't know since it's the only thing written in my homework, ,maybe its any line which pass through this point which is (h,k)

OpenStudy (foolaroundmath):

(y-k)=m*(x-h) vary m, which is the slope of the line, you'll get a new line for each value of m

OpenStudy (anonymous):

Nevermind, I see what it meant. You need to generate a line with the point-slope form, \[y - y_1 = m(x-x_1)\] Now, you can set y_1 = k and x_1 = h. Change m in any way you like, and you'll generate all lines (in effect, rotated about (h,k) that pass through (h,k).

OpenStudy (anonymous):

Thank you:))